OpenStudy (anonymous):

Find the indicated probability, given that the lifetime of a lightbulb is exponentially distributed with pdf f(x) = 6e^-6x (with x measured in years). The lightbulb lasts less than 3 months.

OpenStudy (anonymous):

What class is this for?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

@rizwan_uet can u help?

OpenStudy (anonymous):

i m very weak at probability sorrr @onegirl

OpenStudy (anonymous):

*sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@cwrw238 can u help?

OpenStudy (zarkon):

convert 3 months to years and integrate

OpenStudy (zarkon):

\[P(X<a)=P(0\le X<a)=\int\limits_{0}^{a}6e^{-6x}dx\]

OpenStudy (zarkon):

where \(a\) is in years

OpenStudy (anonymous):

so i get 0.25 so whats next?

OpenStudy (anonymous):

so i will substitute 0.25 into a and solve.

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

okay thanks